Prove that is a body is thrown vertically upwards , then the time of ascent is equal to the time of descent .

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT
Suppose a body is projected in the upward direction with a velocity u and H is maximum height attained by the body.
The velocity of the body at maximum height will be zero, i.e. v = 0.
Acceleration due to gravity (g) acts on body in the downward direction.
From the first equation of motion, 
v=u+at
0=ugtascent
tascent=u/g

Let t be total time of flight. From the second equation of motion,
S=ut+1/2at2
0=t(u1/2gt)
t=0,t=2ug
Now, tascent+tdescent=t=2u
/g
tdescent=u/g

tascent=tdescent=u/g